# Twisting Lids Off with Two Hands

### v1

@@ Abstract
Manipulating objects with two multi-fingered hands has been a long-standing challenge in robotics, attributed to the contact-rich nature of many manipulation tasks and the complexity inherent in coordinating a high-dimensional bimanual system. In this work, we consider the problem of twisting lids of various bottle-like objects with two hands, and demonstrate that policies trained in simulation using deep reinforcement learning can be effectively transferred to the real world. With novel engineering insights into physical modeling, real-time perception, and reward design, the policy demonstrates generalization capabilities across a diverse set of unseen objects, showcasing dynamic and dexterous behaviors. Our findings serve as compelling evidence that deep reinforcement learning combined with sim-to-real transfer remains a promising approach for addressing manipulation problems of unprecedented complexity.
{NoHyper}* Equal Contribution.
I Introduction
@@
We study the task of twisting or removing lids with two multi-fingered robot hands. This task is both practically important and profoundly interesting. For one, the ability to twist or remove lids from containers is a crucial motor skill that toddlers acquire during their early developmental stages [36, 54]. For another, the manipulation skills required for this task, such as the coordination of fingers to manipulate a multi-part object, can be generally useful across a large collection of practical tasks.
Our work demonstrates the feasibility of learning a dexterous and dynamic bimanual manipulation policy purely in simulation and zero-shot transferring it to the real world. By abstracting a class of articulated objects into a simplistic two-part object model with randomized physical properties, we successfully train a policy that can generalize to a diverse set of novel objects with different shapes, sizes, visual appearances, and other physical properties (Figure LABEL:fig:teaser). Our method does not require precise modeling of any individual object, or hardcoding prior knowledge on object properties. Instead, it allows for stable and natural twisting behaviors to emerge through large-scale reinforcement learning (RL) training.
In developing a system with these capabilities, we uncover a number of novel engineering insights, which we outline below.
@@
Physical Modeling. Our work features a novel class of objects for in-hand manipulation: articulated objects defined as two rigid bodies connected via a revolute joint with threaded structure. Accurately modeling friction and contact with revolute joints and threaded structure has traditionally been a hard challenge in robotics simulation [33]. To address this, we introduce a brake-based design to model the interaction between lid and body of bottle-like objects. This design is fast to simulate while maintaining a high fidelity to real-world physical dynamics, enabling efficient policy learning and successful sim-to-real transfer.
@@
Perception. We initially hypothesize that a fine-grained, contact-rich manipulation task like lid-twisting must require precise perceptual information on object states. To our surprise, a two-point sparse object representation, extracted from off-the-shelf object segmentation and tracking tools, is sufficient to solve the perception problem. With simple domain randomization techniques, we train policies that are robust against occlusion and camera noise. This discovery suggests that a minimal amount of object perception information can be adequate for complicated bimanual manipulation tasks.
@@
Reward Design. Previously performant reward designs for tasks like in-hand reorientation cannot be straightforwardly applied to our task, since those tasks focus on manipulating single-part rigid bodies with one hand rather than multi-part articulated bodies with two hands. Solving this task is more challenging since it involves more complex and precise contact (e.g. using two hands to hold a lid). In addressing this challenge, we discover a simple keypoint-based contact reward that yields natural lid-twisting behavior on the robot fingers.
@@
We conduct several controlled experiments in both simulation and the real world. Through empirical analysis, we verify that our simulation modeling, perception module, and reward design can reliably lead to the desired behavior of lid twisting. Our final successful policy manifests natural behavior across test objects with various physical properties such as shapes, sizes, and masses in simulation. Moreover, the learned policy can be zero-shot transferred to a wide range of novel household objects whose lids can be removed (Figure LABEL:fig:teaser), and it is robust against perturbations.
II Background
II-A Classic Approaches for Bimanual Manipulation
For decades, the problem of bimanual manipulation has remained a fascinating unsolved challenge in robotics [35, 23, 48, 53, 6, 49]. While multi-fingered robot hands seem to be a natural choice for bimanual robot systems in theory, designing controllers for high-dimensional action spaces remains an open problem due to difficulties in mechanical design and reliable actuation. Most previous works in bimanual manipulation thus use simple and durable parallel jaw-grippers as end-effectors. For example, Caccavale et al. [5] demonstrate carrying and moving objects using two arms with impedance control. Sarkar et al. [45] show controlling object states using rolling contact. Unlike most previous works, our work focuses on bimanual manipulation with two multi-fingered hands.
While there exist prior works that utilize two multi-fingered hand for manipulation, they differ substantially from our work in terms of approach and results. Vahrenkamp et al. [53] presents a bimanual grasp planner that could be used to build grasps for large objects using multi-fingered hands, but without demonstrating any real-world results. Platt et al. [39] achieve object reconfiguration using two three-fingered robot hands, but do so by planning a sequence of low-level controllers. Steffen et al. [50] study a screwing task similar to our task of interest but synthesize the control motion sequence with a kernel approach applied to human motion data captured with a data glove. Our work fills a missing space in the bimanual multi-fingered manipulation literature, as we do not rely on any planner, accurate object model, or human data. Instead, we directly work on a high-dimensional action space with RL on randomized environments.
II-B Learning Approaches for Bimanual Manipulation
@@
In recent years, bimanual manipulation has been more actively studied with learning-based methods, as a result of progress in learning algorithms and compute infrastructure. These learning-based approaches can generally be categorized into two types: 1) learning from real-world data; 2) learning in simulation, then transferring to the real world (sim-to-real).
@@
Learning from Real-World Data. Rapid progress has been made in reinforcement learning (RL) in the real world. Zhang et al. [58] learn to chain motor primitives for vegetable cutting, with relatively motion primitives; much of the task difficulty is bypassed via the use of specialized end-effectors [1, 14]. Chiu et al. [12] learn precise needle manipulation with two grippers by integrating RL with a sampling-based planner. While impressive, these works cannot easily scale to higher dimensional action space due to their sample inefficiency or the need to define heuristic-based action primitives.
@@
Most recent successes in bimanual manipulation are achieved by learning from demonstrations [47, 51, 59]. However, successes so far are largely limited to simple end-effectors like parallel jaw grippers. One reason is the lack of high-quality demonstration data from multi-fingered robot hands [24]; the availability of such hand hardware is itself limited, not to mention the sophisticated data collection infrastructure. Although several works aiming to improve demonstration data collection with two arms [28, 26, 15] or multi-fingered hands exist [2, 3, 42, 43, 16], their latency and retargeting errors limit their practical applicability and scalability. Our method uses RL in simulation and is thus not limited by the hardware and data collection infrastructure problems faced by learning from demonstration approaches.
@@
Sim-to-Real. There has been growing interest in sim-to-real approaches for robotics – i.e. learning policies in simulation and transferring them to the real world – stimulated by several notable successes in recent years ranging from locomotion [19, 32, 25] to manipulation [8, 17, 34, 40]. Existing works in manipulation, however, are mostly done with either a single multi-fingered hand [7, 21, 38, 41, 44, 55, 56, 10, 52], or two arms with simpler end-effectors [30, 27, 20]. While Chen et al. [9] and Zakka et al. [57] feature bimanual tasks with dexterous hands, only simulation results are shown. Perhaps the work most related to ours is Huang et al. [18], where the authors demonstrate throwing and catching objects using two dexterous hands. However, our task is significantly more contact-rich and requires substantially more challenging bimanual coordination to maintain object stability at all times. Our work distinguishes itself from the literature by considering manipulation for an extremely contact-rich task using two multi-fingered hands; notably, the objects that we work with are two-part articulated bodies, while previous work focuses on single-part rigid bodies.
Figure 1: Bottle URDF model. A key design in the bottle model is the brake link and the prismatic “brake” joint. The brake link constantly presses against the lid link to simulate the static friction between the base link and the lid link. Without the brake, the lid link can easily start to rotate even when it is not in contact with the fingers.
III Task Formulation
Twisting lids of container objects is a complex in-hand manipulation process that requires dynamic dexterity of multiple fingers and precise coordination between two hands. Below, we define the specific task being considered in this work.
Task Initialization. Each object of interest consists of two rigid, near-cylindrical parts (a “body” and a “lid”); the two parts are connected via a continuous revolute joint, allowing them to rotate about each other. To better benchmark the bimanual twisting capability, we consider a class of articulated bottles with lids that can be twisted infinitely (see Section IV-B for more details). The two robotic hands are initialized in a static pose with upward-facing palms. At the beginning of each episode, a bottle-like object is gently dropped or placed onto the fingers. The initial pose of the object is randomized both in translation and rotation to a fixed default pose; the initial joint positions of the hands are randomized about a predefined canonical pose by adding Gaussian noise. Note that since we do not assume a stable grasp configuration at task initialization, the control policy needs to learn in-grasp reorientation to place the object in a stable location to perform successive manipulation.
@@ III Task Formulation
Task Objective. The goal of this task is to twist the lid about the object’s axis of rotation in one direction as much as possible; during this process, the object should always stay in hand. Achieving this involves a sequence of delicate movements: 1) after initialization, the robot hand should firmly grasp and slightly rotate the bottle to a suitable pose; 2) the hand that is closer to the object lid should place its finger around the lid to initiate twisting motion; 3) the two hands should coordinate to avoid dropping the object while one hand twists the lid. Motor skills that arise from this task could serve as generic abstractions for skills necessary to manipulate many other household objects, especially those with revolute joints such as Rubik’s cubes, light bulbs, and jars.
Task Complexity. The complexity inherent in this task can be summarized into the following three aspects.
1) In-Hand Perception Difficulty. Heavy occlusion can occur during manipulation due to fingers surrounding the object in-hand. How to build a real robot system that can perform robust bimanual manipulation against such occlusion and perceptual noises remains an open problem.
2) Fine-grained Contact. This task requires both hands to coordinate to maintain stability. Incorrect contact points or forces could result in the bottle dropping.
3) High-Dimensional Control. Our system has $32$ degrees of freedom, and almost all the fingers are engaged throughout the manipulation process. This requires extremely precise coordination among a large number of rigid bodies, which is a known challenge for existing learning-based methods.
IV System Setup
To address the aforementioned challenges in bimanual manipulation and our specific task choice, we develop a sim-to-real pipeline based on a deep RL approach. In this section, we describe our overall system setup.
IV-A Real-world System
@@
Figure 2: Real-time perception system. Top: scene overview. Bottom: we segment and track object parts from the RGB frames (left), take mask centers as object part centers (middle), and estimate 3D object keypoints using noisy depth information from the camera (right).

### v2

@@ Abstract
Manipulating objects with two multi-fingered hands has been a long-standing challenge in robotics, due to the contact-rich nature of many manipulation tasks and the complexity inherent in coordinating a high-dimensional bimanual system. In this work, we share novel insights into physical modeling, real-time perception, and reward design that enable policies trained in simulation using deep reinforcement learning (RL) to be effectively and efficiently transferred to the real world. Specifically, we consider the problem of twisting lids of various bottle-like objects with two hands, demonstrating policies with generalization capabilities across a diverse set of unseen objects as well as dynamic and dexterous behaviors. To the best of our knowledge, this is the first sim-to-real RL system that enables such capabilities on bimanual multi-fingered hands.
***Correspondence to {toru,zhaohengyin}@berkeley.edu. The first two authors contribute equally.
Figure 1: We train two anthropomorphic robot hands to twist (off) lids of various articulated objects. The control policy is first trained in simulation with deep reinforcement learning, then zero-shot transferred to a real-world setup. We show that a single policy trained to manipulate simplistic, simulated bottle-like objects can generalize to real-world objects that have drastically different physical properties (e.g. shape, size, color, material, mass). The length, diameter (or diagonal length), and mass of each object are annotated at the bottom of individual subfigures. More results can be found in our video and and project website.
Keywords: Bimanual Manipulation, Sim-to-Real, Reinforcement Learning
1 Introduction
@@
Achieving dexterous bimanual manipulation with two anthropomorphic robot hands has been exceptionally challenging, due to the inherently contact-rich nature of many manipulation tasks and the complexity of coordinating a high-dimensional bimanual system. This work takes a step towards this grand goal, by demonstrating the feasibility of learning a highly dexterous and dynamic bimanual manipulation policy purely in simulation and zero-shot transferring it to the real world. Specifically, we study the task of twisting or removing lids with two multi-fingered robot hands. This task is both practically important and profoundly interesting: for one, the ability to twist or remove lids from containers is a crucial motor skill that toddlers acquire during their early developmental stages [1, 2]; for another, the manipulation skills required for this task, such as the coordination of fingers to manipulate a multi-part object, can be generally useful across a large collection of practical tasks.
Since collecting human expert demonstration data to solve contact-rich tasks via imitation learning is highly challenging and expensive [3], we aim at training a generalizable policy through sim-to-real reinforcement learning without using any expert data (Figure 1). Our method does not require precise modeling of any individual object, or hardcoding prior knowledge on object properties; instead, stable and natural bimanual finger behaviors emerge through large-scale reinforcement learning (RL) training. Below, we share the novel insights that enable us to develop such a system.
@@
Physical Modeling. Our work features a novel class of objects for in-hand manipulation: articulated objects defined as two rigid bodies connected via a revolute joint with a threaded structure. Accurately modeling friction and contact with revolute joints and threaded structures has traditionally been a hard challenge in robotic simulation [4]. To address this, we introduce a brake-based design to model the interaction between the lid and body of bottle-like objects. This design is fast to simulate while maintaining high fidelity to real-world physical dynamics, enabling efficient policy learning and successful sim-to-real transfer.
@@
Perception. We initially hypothesize that a fine-grained, contact-rich manipulation task like lid-twisting must require precise perceptual information on object states and shapes. To our surprise, a two-point sparse object representation, extracted from off-the-shelf object segmentation and tracking tools, is sufficient to solve the perception problem. With simple domain randomization techniques, we train policies that are robust against occlusion and camera noise. This discovery suggests that a minimal amount of perception information can be adequate for complicated bimanual manipulation tasks.
@@
Reward Design. Previously performant reward designs for tasks like in-hand reorientation [5, 6, 7] cannot be straightforwardly applied to our task, since those tasks focus on manipulating single-part rigid bodies with one hand rather than multi-part articulated bodies with two hands. Solving this task is more challenging since it involves more complex and precise contact (e.g. using two hands to hold a lid). In addressing this challenge, we discover a simple keypoint-based contact reward that yields natural lid-twisting behavior on the robot fingers.
@@
We conduct several controlled experiments in both simulation and the real world. Through empirical analysis, we verify that our simulation modeling, perception module, and reward design can reliably lead to the desired behavior of lid-twisting. Our final successful policy manifests natural behavior across test objects with various physical properties such as shapes, sizes, and masses in simulation. Moreover, the learned policy can be zero-shot transferred to a wide range of novel household objects whose lids can be removed (Figure 1), and it is robust against perturbations.
2 Background
@@
For decades, bimanual manipulation has remained an unsolved challenge in robotics [8, 9, 10, 11, 12, 13]. While multi-fingered robot hands seem to be a natural choice for bimanual robot systems in theory, designing controllers for high-dimensional action spaces remains an open problem. Classical approach has made significant progress but usually assume known object or physics information [14, 15] and the generalizability remains unknown. In recent years, bimanual manipulation has been actively studied with learning-based methods, as a result of progress in learning algorithms and compute infrastructure. These learning-based approaches can be categorized into two types: 1) learning from real-world data; 2) learning in simulation, then transferring to the real world (sim-to-real).
@@
Learning from Real-World Data. Rapid progress has been made in RL in the real world. Zhang et al. [16] learn to chain motor primitives for vegetable cutting, with relatively simple motion primitives; much of the task difficulty is bypassed via the use of specialized end-effectors [17, 18]. Chiu et al. [19] learn precise needle manipulation with two grippers by integrating RL with a sampling-based planner. While impressive, these works cannot easily scale to higher dimensional action space due to their sample inefficiency or the need to define heuristic-based action primitives.
@@
Most recent successes in bimanual manipulation are achieved by learning from demonstrations [20, 21, 22, 23, 24, 25]. However, successes so far are largely limited to simple end-effectors like parallel jaw grippers due to the lack of high-quality demonstration data from multi-fingered robot hands [26]. Although several works aiming to improve demonstration data collection with two arms [27, 28, 29] or multi-fingered hands [21, 30, 31, 32, 33], their latency and retargeting errors limit their practical applicability and scalability. Lin et al. [20] proposes a scalable hands-arms teleoperation system that learns smooth bimanual policies, but the system compromises on dexterity. Similar systems that offer more dexterous control [34, 35], on the other hand, suffer from drawbacks ranging from jittery control to high costs. Our method uses RL in simulation and is thus not limited by the hardware and data collection infrastructure problems faced by learning from demonstration approaches.
@@
Sim-to-Real. There has been growing interest in sim-to-real approaches for robotics – i.e. learning policies in simulation and transferring them to the real world – stimulated by several notable successes in recent years ranging from locomotion [36, 37, 38] to manipulation [5, 7, 6, 39]. Existing works in manipulation, however, are mostly done with either a single multi-fingered hand [3, 40, 41, 42, 43, 44, 45, 46, 47], or two arms with simpler end-effectors [48, 49, 50]. While Chen et al. [51] and Zakka et al. [52] feature bimanual tasks with dexterous hands, only simulation results are shown. The work most related to ours is Huang et al. [53], where the authors demonstrate throwing and catching objects using two dexterous hands. However, our task is significantly more contact-rich and requires substantially more challenging bimanual coordination to maintain object stability at all times. To our best knowledge, there is no learning-based method directly comparable to ours on the proposed task.
Figure 2: Our bottle model and the used bottles in the simulation and the real world. A: Simulated bottle URDF. B: Training bottle objects in simulation. C: Custom-made bottles (in-distribution except for the rightmost square bottle). D: Household object bottles (out-of-distribution).
3 Learning to Twist Lids with Two Hands
@@
We focus on the challenging task of lid-twisting for container objects, since it is a complex in-hand manipulation process that requires dynamic dexterity of multiple fingers and precise coordination between two hands. The goal of this task is to twist the lid about the object’s axis of rotation in one direction as much as possible; during this process, the object should always stay in hand. Achieving this involves a sequence of delicate movements: 1) after initialization, the robot hand should firmly grasp and slightly rotate the bottle to a suitable pose; 2) the hand that is closer to the object lid should place its finger around the lid to initiate twisting motion; 3) the two hands should coordinate to avoid dropping the object while one hand twists the lid. Motor skills that arise from this task could serve as generic abstractions for skills necessary to manipulate many other household objects, especially those with revolute joints such as Rubik’s cubes, light bulbs, and jars.
3.1 Object Simulation
A central challenge in simulating the lid-twisting task is how to model friction between the bottle body and the lid properly, particularly static friction. Simulating this type of physical force has been a long-standing problem in robotics and graphics [4]. We design a simple modeling approximation that strikes a balance between fidelity and speed during physical simulation; our bottle-like object model is illustrated in Figure 2(A). Our design features a special Brake Link that constantly presses against the bottle lid via a prismatic joint. This artificially generates frictional forces between the bottle body (Base Link) and the lid (Lid Link), preventing relative rotation between them — similar to a bottle with its lid screwed on. We replicate these bottles in the real world, as shown in Figure 2(B).
We note that naively tuning static friction properties between two revolute-joint-connected bodies is not realistic enough with our simulator. Such a brake-based is the only way we find that can simulate the static friction well.
3.2 Task Initialization
To better benchmark the bimanual twisting capability, we consider a class of articulated bottles with lids that can be twisted infinitely (see Figure 2(A) and Section 3.1 for more details). Each object of interest consists of two rigid, near-cylindrical parts (a “body” and a “lid”); the two parts are connected via a continuous revolute joint, allowing them to rotate about each other. At the beginning of each episode, the two robotic hands are initialized in a static pose with upward-facing palms, and a bottle-like object is gently dropped or placed onto the fingers. The initial pose of the object is randomized both in translation and rotation to a fixed default pose; the initial joint positions of the hands are randomized about a canonical pose by adding Gaussian noise. Note that since we do not assume a stable grasp configuration at task initialization, the control policy needs to learn in-grasp reorientation to place the object in a stable location to perform successive manipulation.
Figure 3: Left: Real-time perception system. Top: overview. Bottom: we segment and track object parts from the RGB frames (left), take mask centers as object part centers (middle), and estimate 3D object keypoints using noisy depth information from the camera (right). Right: Illustration of reward design. Our task-specific reward contains finger contact reward (yellow arrows), twisting reward (white arrow), and pose reward (blue arrow). In particular, our keypoint-based finger contact reward is crucial for learning the desired behavior.
3.3 Policy Learning
Bimanual in-hand dexterous manipulation involves highly complex hand-object contacts, and remains challenging to solve with traditional methods. In this work, we address the control challenge through RL. We formulate our control problem as a partially observable Markov Decision Process.
Action Space. We use a PD controller to drive the robot hand. The control policy produces a relative target joint position as the action $a_{t}$, which is added to the current target joint position $q_{t}$ to produce the next target: $\tilde{q}_{t+1}=\tilde{q}_{t}+\eta{\rm EMA}(a_{t})$. $\eta$ is a scaling factor. Note that we smooth the action with its exponential moving average (EMA) to produce smooth motion. The next target position is sent to the PD controller to generate torque on each joint.
Reward. While one way to approach hard exploration problems is to add intrinsic rewards [54, 55], we introduce the following fine-grained reward terms to shape the hand behavior (Figure 3 right). (1). Twisting Reward. We define the twisting reward as $r_{\mathrm{twisting}}=\Delta\theta=q^{t+1}_{bottle}-q^{t}_{bottle}$ which is the rotation angle of the lid during one-step execution. This reward term encourages the hand to twist the lid. (2). Finger Contact Reward. We find it crucial to use a set of reference contact points to guide effective contact between fingertips and the bottle. We define two set of points $\mathbf{X}^{L}\in\mathbb{R}^{n\times 3}$ and $\mathbf{X}^{R}\in\mathbb{R}^{m\times 3}$ attached on the bottle base and lid respectively. Then, we define the finger contact reward as $r_{\mathrm{contact}}=\sum_{i}\left[\frac{1}{1+\alpha d(\mathbf{X}^{L},\mathbf{F}^{L}_{i})}+\frac{1}{1+\alpha d(\mathbf{X}^{R},\mathbf{F}^{R}_{i})}\right],$ where $\mathbf{F}^{L}\in\mathbb{R}^{4\times 3}$ and $\mathbf{F}^{R}\in\mathbb{R}^{4\times 3}$ are the position of left and right fingertips, $\alpha$ is a scaling hyperparameter, and $d$ is a distance function defined as $d(\mathbf{A},\mathbf{x})=\min_{i}\|\mathbf{A}_{i}-\mathbf{x}\|_{2}.$ Therefore, we require each fingertip to stay as close to one of the reference contact points as possible. As we will see later, this term is necessary for eliciting desired behavior and task success.
(3). Pose Reward. We also introduce a pose matching reward term to encourage the bottle main axis $\mathbf{x}_{axis}$ aligned with a predefined direction $\mathbf{v}$. This term is defined as $r_{\mathrm{pose}}=-\arccos(\langle\mathbf{x}_{axis},\mathbf{v}\rangle).$ (4). Regularizations. Besides the three task-specific rewards, we also introduce another few regularization terms as in previous works [44], including work penalty and action penalty to penalize large, jerky motions. We leave the details of the definition to the appendix.
Reset Strategy. There exist many possible hand-object interaction modes. Among these, most modes lead to failures such as getting the object stuck between fingers, and exploring those modes rarely provides good learning signals. To circumvent the high dimensionality of our exploration problem, we introduce two early termination criteria. First, we reset an episode if the robot hands fail to rotate the bottle into a desired pose for bimanual twisting within a short time limit. Additionally, we reset when the bottle’s $z$-position is below a certain threshold, as the fingertips of the two hands can pinch the bottle at a low position without being able to reposition it into the palm.
Training. We use PPO [56] with asymmetric critic observation [57] to train our policy, and introduce various domain randomizations to make the policy transferable to the real world. We apply both both physical and non-physical randomizations. The detailed training setup can be found in the appendix.
Figure 4: Training curves in different settings. Top: Single-object training results (evaluated on single-object setup). Bottom: Multi-object training results (evaluated on multi-object setup). Left half: Comparisons of different reward setups. Right half: Ablations on the use of vision. The results are averaged on 5 seeds. The shaded area shows the standard deviation. The AD score is averaged by the total execution steps.
3.4 Real-World Perception.
@@ 3.3 Policy Learning
4 Simulation Experiments
To test how to enable the emergence of natural and robust manipulation behaviors, we first conduct several experiments in simulation. Specifically, we study the following questions: How important is the keypoint-based reward for eliciting desired twisting behavior in bimanual manipulation? How important is visual information for solving this task? Is a sparse keypoint representation enough for learning a generalizable policy?
4.1 Setup
Object Set. In simulated experiments, we use a collection of simulated cylindrical bottles with varying aspect ratios for both training and evaluation. Some samples are visualized in Figure 2(B). We consider two setups in simulation: 1) multi-object, in which all the objects are used, and 2) single-object, in which a single medium-sized bottle that represents the mean of the dataset is used.
Evaluation Metric. We introduce the following metrics for evaluating the performance: 1) Angular Displacement (AD) is the total number of degrees through which the lid has been twisted; 2) Time-to-Fail (TTF) is the period measured from the moment the bottle is held to the point when it either slips from the hand or becomes lodged; 3) Velocity (Vel) is AD divided by TTF, reflecting the speed of twisting motion.
Baselines. We compare our policy with the following baselines. 1) Policy without Vision. This is a neural network policy without object state information. We use this to evaluate the importance of vision. 2) Policy with Reduced Contact Reward. In training this policy, we reduce the intensity of our proposed finger contact reward. We use this to study the role of our contact reward in policy learning and shaping the policy’s behavior. 3) Policy with Gait Constraint Reward. In training this policy, we replace our contact reward with a gait constraint reward function similar to ones used for in-hand reorientation tasks [39]. This baseline is only used for qualitative analysis.
Figure 5: Left: Behavior of different reward functions. Top: Our full reward function achieves a stable grasp, as well as a smooth, natural, and human-like twisting motion. Middle: A naive gait constraint reward without any contact hints leads to erratic finger motion and unnatural grasps. Bottom: A reduced contact reward yields somewhat natural behavior, but the grasp is loose compared to the full contact reward case. Right: Perturbing a learned policy with random external force. Our policy is resilient to these external forces and able to recover.
4.2 Results
Reward Design. We first compare our approach with the reduced finger reward baseline (Figure 4). After decreasing the scale of finger contact reward, learned policies fail to master the desired lid-twisting skill and have low performance in general. We hypothesize that this is because the motion of lid-twisting requires a very specific pose pattern for holding the object; without explicitly encouraging such a pose pattern (e.g., via its contact modes), RL exploration becomes so hard that it is unsolvable within the available training time. We also observe a positive correlation between the intensity of finger contact reward and both 1) sample efficiency during learning and 2) performance of learned policies (as reflected by Figure 4 and qualitative observations in Figure 5 (left)).
Vision vs. No Vision. We also study the importance of vision modality. Existing works show that certain rotation behaviors can be achieved through implicit tactile sensing (via proprioception) [39]. However, our empirical results show that, in both single and multi-object setups, the no-vision baseline performs substantially worse than our full method. This suggests that knowledge of the position of bottle keypoints is essential for successful lid-twisting.
Single Object vs Multi Object. We run RL training with two object settings: 1) using a single bottle-like object; 2) using multiple bottle-like objects with more variation in the ratio between the bottle base and lid. For results shown in Figure 4, all multi-object training runs are evaluated on multi-object setup and all single-object training runs are evaluated on single-object setup. The two settings pose a trade-off between specialization and generalization: in the single-object scenario, the policy might learn successful behaviors more easily but find it harder to generalize to unseen objects, and vice versa. To our surprise, we observe that multi-object training yields slightly better performance compared to single-object training. We hypothesize that multi-object makes exploring lid-twisting behavior an easier process by introducing an object curriculum that covers both easy and hard object instances during training.
Table 1: Comparison with baselines on real setup. For each method, we deploy 3 policies trained on 3 different seeds and average the results. Each deployment trial is conducted for 30 seconds.
BlueBottle WoodBottle RedBottle GoldBottle SqaureBottle
Method AD $\uparrow$ TTF $\uparrow$ Vel $\uparrow$ AD $\uparrow$ TTF $\uparrow$ Vel $\uparrow$ AD $\uparrow$ TTF $\uparrow$ Vel $\uparrow$ AD $\uparrow$ TTF $\uparrow$ Vel $\uparrow$ AD $\uparrow$ TTF $\uparrow$ Vel $\uparrow$
Replay $\mathop{\ 128.33}\limits_{\pm 217.96}$ $\mathop{\ 7.67}\limits_{\pm 4.93}$ $\mathop{\ 11.68}\limits_{\pm 19.80}$ $\mathop{\ 2.67}\limits_{\pm 4.62}$ $\mathop{\ 7.67}\limits_{\pm 5.86}$ $\mathop{\ {0.22}}\limits_{\pm{0.38}}$ $\mathop{\ 15.00}\limits_{\pm 25.98}$ $\mathop{\ 4.67}\limits_{\pm 4.62}$ $\mathop{\ {1.59}}\limits_{\pm{2.60}}$ $\mathop{\ 28.33}\limits_{\pm 43.04}$ $\mathop{\ 7.67}\limits_{\pm 4.04}$ $\mathop{\ {2.99}}\limits_{\pm{4.17}}$ $\mathop{\ 29.67}\limits_{\pm 8.62}$ $\mathop{\ 10.00}\limits_{\pm 0.00}$ $\mathop{\ {2.97}}\limits_{\pm{0.86}}$
No-Vis $\mathop{\ 1.33}\limits_{\pm 2.31}$ $\mathop{\ 21.67}\limits_{\pm 14.43}$ $\mathop{\ {0.04}}\limits_{\pm{0.08}}$ $\mathop{\ 1.07}\limits_{\pm 1.85}$ $\mathop{\ 14.67}\limits_{\pm 13.61}$ $\mathop{\ {0.27}}\limits_{\pm{0.46}}$ $\mathop{\ 1.90}\limits_{\pm 3.29}$ $\mathop{\ 8.33}\limits_{\pm 6.11}$ $\mathop{\ {0.27}}\limits_{\pm{0.47}}$ $\mathop{\ 0.67}\limits_{\pm 1.15}$ $\mathop{\ 16.33}\limits_{\pm 13.05}$ $\mathop{\ {0.04}}\limits_{\pm{0.08}}$ $\mathop{\ 5.00}\limits_{\pm 6.24}$ $\mathop{\ 20.33}\limits_{\pm 11.24}$ $\mathop{\ {0.18}}\limits_{\pm{0.20}}$
No-Asym $\mathop{\ 18.67}\limits_{\pm 28.94}$ $\mathop{\ 30.00}\limits_{\pm 0.00}$ $\mathop{\ {0.62}}\limits_{\pm{0.96}}$ $\mathop{\ 0.67}\limits_{\pm 1.15}$ $\mathop{\ 19.33}\limits_{\pm 15.14}$ $\mathop{\ {0.03}}\limits_{\pm{0.04}}$ $\mathop{\ 8.33}\limits_{\pm 14.43}$ $\mathop{\ 13.00}\limits_{\pm 15.13}$ $\mathop{\ {0.28}}\limits_{\pm{0.48}}$ $\mathop{\ 4.3}\limits_{\pm 7.51}$ $\mathop{\ 3.67}\limits_{\pm 3.79}$ $\mathop{\ {0.54}}\limits_{\pm{0.94}}$ $\mathop{\ 0.00}\limits_{\pm 0.00}$ $\mathop{\ 2.33}\limits_{\pm 1.53}$ $\mathop{\ {0.00}}\limits_{\pm{0.00}}$
Large $\mathop{\ 2.00}\limits_{\pm 2.00}$ $\mathop{\ 22.33}\limits_{\pm 13.28}$ $\mathop{\ {0.14}}\limits_{\pm{0.14}}$ $\mathop{\ 0.00}\limits_{\pm 0.00}$ $\mathop{\ 24.00}\limits_{\pm 10.39}$ $\mathop{\ {0.00}}\limits_{\pm{0.00}}$ $\mathop{\ 2.33}\limits_{\pm 2.52}$ $\mathop{\ 9.33}\limits_{\pm 4.73}$ $\mathop{\ {0.47}}\limits_{\pm{0.68}}$ $\mathop{\ 2.67}\limits_{\pm 3.06}$ $\mathop{\ 22.33}\limits_{\pm 13.28}$ $\mathop{\ {0.09}}\limits_{\pm{0.10}}$ $\mathop{\ 1.67}\limits_{\pm 2.89}$ $\mathop{\ 30.00}\limits_{\pm 0.00}$ $\mathop{\ {0.06}}\limits_{\pm{0.10}}$
Ours $\mathop{\ \textbf{946.33}}\limits_{\pm\textbf{383.81}}$ $\mathop{\ \textbf{23.67}}\limits_{\pm\textbf{10.97}}$ $\mathop{\ \textbf{41.26}}\limits_{\pm\textbf{4.36}}$ $\mathop{\ \textbf{499.50}}\limits_{\pm\textbf{578.23}}$ $\mathop{\ \textbf{30.0}}\limits_{\pm\textbf{0.0}}$ $\mathop{\ \textbf{16.65}}\limits_{\pm\textbf{19.27}}$ $\mathop{\ \textbf{150.67}}\limits_{\pm\textbf{113.47}}$ $\mathop{\ \textbf{30.00}}\limits_{\pm\textbf{0.00}}$ $\mathop{\ \textbf{5.02}}\limits_{\pm\textbf{3.78}}$ $\mathop{\ \textbf{98.67}}\limits_{\pm\textbf{66.91}}$ $\mathop{\ \textbf{30.00}}\limits_{\pm\textbf{0.00}}$ $\mathop{\ \textbf{3.29}}\limits_{\pm\textbf{2.23}}$ $\mathop{\ \textbf{43.00}}\limits_{\pm\textbf{12.12}}$ $\mathop{\ \textbf{30.00}}\limits_{\pm\textbf{0.00}}$ $\mathop{\ \textbf{1.43}}\limits_{\pm\textbf{0.40}}$
5 Real-world Experiments
5.1 Experiment Setup